A 20.6-L tank contains 0.206 kg of helium at 16.4°C. The molar mass of helium is 4.00 g/mol.

(a) How many moles of helium are in the tank?
  mol

(b) What is the pressure in the tank, in pascals and in atmospheres?

Pa ?
atm?
Answer #1

a) Number of moles n=\frac{206}{4}=51.5moles .\implies P=\frac{51.5\times8.31\times289.4}{20.6\times10^{-3}}

b)PV=nRT

\implies P=\frac{nRT}{V}

\implies P=\frac{51.5\times8.31\times289.4}{20.6\times10^{-3}}, All in SI units

\implies P=6012285 Pa=59.33atm